\begin{thebibliography}{TukV2}




\bibitem [Ahl1] {Ahlfors1} L.~Ahlfors, {\it Complex Analysis},
McGraw-Hill, 1979.

\bibitem [Ahl2] {Ahlfors2} L.~Ahlfors, {\it Lectures on Quasiconformal
Mappings}, Van Nostrand, 1966.

\bibitem [Ahl3] {Ahlfors3} L.~Ahlfors, {\it Conformal Invariants:
Topics in Geometric Function Theory}, McGraw-Hill, 1973.

\bibitem [Bal1] {Ball1} J.~Ball, {\it Convexity conditions and existence
theorems in nonlinear elasticity}, Archive for Rational Mechanics and
Analysis {\bf 63} (1977), 337--403.

\bibitem [Bal2] {Ball2} J.~Ball, {\it Global invertibility of Sobolev
functions and the interpenetration of matter}, Proceedings of the
Royal Society of Edinburgh {\bf 88A} (1981), 315--328.

\bibitem [Bin] {Bing} R.~Bing, {\it The Geometric Topology of
$3$-Manifolds}, Colloquium Publications {\bf 40}, American
Mathematical Society, 1983.

\bibitem [CanKS] {C-K-S} J.~Cantarella, R.~Kusner, and J.~Sullivan,
{\it On the minimum ropelength of knots and links}, Inventiones
Mathematicae {\bf 150} (2002), 257--286.

\bibitem [Cer] {Cer} A.~{\v C}ernavskii, {\it Local contractibility of the
group of homeomorphisms of a manifold}, Math. Sbornik {\bf 8} (1969),
287--333.

\bibitem [Dav] {Daverman} R.~Daverman, {\it Decompositions of
Manifolds}, Academic Press, 1986.

\bibitem [DonS] {D-S} S.~Donaldson and D.~Sullivan, {\it
Quasiconformal $4$-manifolds}, Acta Mathematica {\bf 163} (1989),
181--252.

\bibitem [EdwK] {E-K} R.~Edwards and R.~Kirby, {\it Deformations of
spaces of embeddings}, Annals of Mathematics (2) {\bf 93} (1971),
63--88.

\bibitem [Fis] {Fisher} G.~Fisher, {\it On the group of all
homeomorphisms of a manifold}, Transactions of the American
Mathematical Society {\bf 97} (1960), 193--212.

\bibitem [Ham] {Hamstrom} M.-E.~Hamstrom, {\it Regular mappings and
the space of homeomorphisms on a $3$-manifold}, Memoirs of the
American Mathematical Society {\bf 40}, 1961.

\bibitem [Hat1] {Hatcher1} A.~Hatcher, {\it Linearization in
$3$-dimensional topology}, Proceedings of the International Congress
of Mathematicians (Helsinki, 1978), 463--468, Academi{\ae} Scientiarum
Fennic{\ae}, 1980.

\bibitem [Hat2] {Hatcher2} A.~Hatcher, {\it A proof of the Smale
conjecture, ${\rm Diff}(S\sp{3})\simeq {\rm O}(4)$ }, Annals of
Mathematics (2) {\bf 117} (1983), 553--607.

\bibitem [Hat3] {Hatcher3} A.~Hatcher, {\it Algebraic Topology},
Cambidge University Press, 2002.

\bibitem [HeiKM] {H-K-M} J.~Heinonen, T.~Kilpel\"ainen, and O.~Martio,
{\it Nonlinear Potential Theory of Degenerate Elliptic Equations},
Oxford University Press, 1993.

\bibitem [Kis1] {Kister1} J.~Kister, {\it Small isotopies in Euclidean
spaces and $3$-manifolds}, Bulletin of the American Mathematical
Society {\bf 65} (1959), 371--373.

\bibitem [Kis2] {Kister2} J.~Kister, {\it Isotopies in $3$-manifolds},
Transactions of the American Mathematical Society {\bf 97} (1960),
213--224.

\bibitem [KrP1] {K-P1} Krantz, S.\ and Parks, H.  (1992).  {\it A
Primer of Real Analytic Functions.}  Birkh\"auser, Boston.

\bibitem [KrP2] {K-P2} Krantz, S.\ and Parks, H.  (1999).  {\it The
Geometry of Domains in Space.}  Birkh\"auser, Boston.

\bibitem [KrP3] {K-P3} Krantz, S.\ and Parks, H.  (2002).  {\it The
Implicit Function Theorem: History, Theory, and Applications.}
Birkh\"auser, Boston.

\bibitem [KusS1] {Ku-S1} R.~Kusner and J.~Sullivan, {\it M\"obius
energies for knots and links, surfaces and submanifolds}, in {\it
Geometric Topology, Proceedings of the Georgia Topology Conference,
1993}, 570--604, American Mathematical Society, 1997.

\bibitem [KusS2] {Ku-S2} R.~Kusner and J.~Sullivan, {\it On distortion
and thickness of knots}, in {\it Topology and Geometry in Polymer
Science}, Springer-Verlag, 1998.

\bibitem [Lau] {Lau} F.~Laudenbach, {\it Les $2$-sph\`eres de
${R}\sp 3$ vues par A.~Hatcher et la conjecture de Smale ${\rm
Diff}(S\sp 3)\sim {\rm O}(4)$}, S\'eminaire Bourbaki 1983/84,
Ast\'erisque {\bf 121--122} (1985), 279--293.

\bibitem [LehV] {Lehto-Virtanen} O.~Lehto and K.~Virtanen, {\it
Quasiconformal Mappings in the Plane}, Springer-Verlag, 1973.

\bibitem [Luu1] {Luu1} J.~Luukkainen, {\it Respectful deformation of
bi-Lipschitz and quasisymmetric embeddings}, Annales Academi{\ae}
Scientiarum Fennic{\ae} Mathematica Ser.~A I Math.\ {\bf 13} (1988),
137--177.

\bibitem [Luu2] {Luu2} J.~Luukkainen, {\it Bi-Lipschitz concordance
implies bi-Lipschitz isotopy}, Monatshefte f\"ur Mathematik {\bf 111}
(1991), 35--46.

\bibitem [Luu3] {Luu3} J.~Luukkainen, {\it Lipschitz and
quasiconformal approximation of homeomorphism pairs}, Topology and its
Applications {\bf 109} (2001), 1--40.

\bibitem [LuuV] {LV} J.~Luukkainen and J.~V\"ais\"al\"a, {\it Elements
of Lipschitz topology}, Annales Academi{\ae} Scientiarum Fennic{\ae}
Mathematica Ser.~A I Math.\ {\bf 3} (1977), 85--122.

\bibitem [Maz] {Mazya} V.~Mazya, {\it Sobolev Spaces},
Springer-Verlag, 1985.

\bibitem [Moi] {Moise} E.~Moise, {\it Geometric Topology in Dimensions
$2$ and $3$}, Springer-Verlag, 1977.

\bibitem [Mor] {Morrey} C.~Morrey, {\it Multiple Integrals in the
Calculus of Variations}, Springer-Verlag, 1966.

\bibitem [RanS] {Randall-Schweitzer} D.~Randall and P.~Schweitzer,
{\it On foliations, concordance spaces, and the Smale conjectures}, in
{\it Differential Topology, Foliations, and Group Actions (Rio de
Janeiro, 1992)}, Contemporary Mathematics {\bf 161}, 235--258,
American Mathematical Society, 1994.

\bibitem [Sem1] {SS1} S.~Semmes, {\it Some topics concerning
homeomorphic parameterizations}, Publicacions Matem\`atiques {\bf 45}
(2001), 3--67.

\bibitem [Sem2] {SS2} S.~Semmes, {\it Real analysis, quantitative
topology, and geometric complexity}, Publicacions Matem\`atiques {\bf
45} (2001), 265--333.

\bibitem [SieS] {LS-DS} L.~Siebenmann and D.~Sullivan, {\it On
complexes that are Lipschitz manifolds}, in {\it Geometric Topology,
Proceedings of the Georgia Topology Conference, 1977}, 503--525,
Academic Press, 1979.

\bibitem [Sul] {Sullivan} D.~Sullivan, {\it Hyperbolic geometry and
homeomorphisms}, in {\it Geometric Topology, Proceedings of the Georgia
Topology Conference, 1977}, 543--555, Academic Press, 1979.

\bibitem [TukV1] {TV1} P.~Tukia and J.~V\"ais\"al\"a, {\it
Quasisymmetric embeddings of metric spaces}, Annales Academi{\ae}
Scientiarum Fennic{\ae} Mathematica Ser.~A I Math.\ {\bf 5} (1980),
97--114.

\bibitem [TukV2] {TV2} P.~Tukia and J.~V\"ais\"al\"a, {\it Lipschitz
and quasiconformal approximation and extension}, Annales Academi{\ae}
Scientiarum Fennic{\ae} Mathematica Ser.~A I Math.\ {\bf 6} (1981),
303--342.

\bibitem [V\"ai1] {V1} J.~V\"ais\"al\"a, {\it Piecewise linear
approximation of lipeomorphisms}, Annales Academi{\ae} Scientiarum
Fennic{\ae} Mathematica Ser.~A I Math.\ {\bf 3} (1977), 377--383.

\bibitem [V\"ai2] {V2} J.~V\"ais\"al\"a, {\it Quasiconformal
concordance}, Monatshefte f\"ur Mathematik {\bf 107} (1989), 155--168.


\end{thebibliography}

\documentclass[12pt,leqno,draft]{article}

\newtheorem{theorem}{Theorem}
\newtheorem{lemma}[theorem]{Lemma}
\newtheorem{proposition}[theorem]{Proposition}
\newtheorem{sublemma}[theorem]{Sublemma}
\newtheorem{definition}[theorem]{Definition}
\newtheorem{corollary}[theorem]{Corollary}
\newtheorem{problem}[theorem]{Problem}
\newtheorem{remark}[theorem]{Remark}
\newtheorem{claim}[theorem]{Claim}
\newtheorem{assumptions}[theorem]{Assumptions}
\newtheorem{examples}[theorem]{Examples}
\newtheorem{question}[theorem]{Question}
\newtheorem{sassumptions}[theorem]{Standing Assumptions}
\newtheorem{sassumption}[theorem]{Standing Assumption}
\newtheorem{conjecture}[theorem]{Conjecture}

\newcommand{\begintheorem}{\addtocounter{equation}{1}\begin{theorem}}
\newcommand{\beginlemma}{\addtocounter{equation}{1}\begin{lemma}}
\newcommand{\beginproposition}{\addtocounter{equation}{1}\begin{proposition}}
\newcommand{\beginsublemma}{\addtocounter{equation}{1}\begin{sublemma}}
\newcommand{\begindefinition}{\addtocounter{equation}{1}\begin{definition}}
\newcommand{\begincorollary}{\addtocounter{equation}{1}\begin{corollary}}
\newcommand{\beginproblem}{\addtocounter{equation}{1}\begin{problem}}
\newcommand{\beginremark}{\addtocounter{equation}{1}\begin{remark}}
\newcommand{\beginclaim}{\addtocounter{equation}{1}\begin{claim}}
\newcommand{\beginassumptions}{\addtocounter{equation}{1}\begin{assumptions}}
\newcommand{\beginexamples}{\addtocounter{equation}{1}\begin{examples}}
\newcommand{\beginquestion}{\addtocounter{equation}{1}\begin{question}}
\newcommand{\beginsassumptions}{\addtocounter{equation}{1}\begin{sassumptions}}
\newcommand{\beginsassumption}{\addtocounter{equation}{1}\begin{sassumption}}
\newcommand{\beginconjecture}{\addtocounter{equation}{1}\begin{conjecture}}

\begin{document}

\title{Some remarks about homeomorphisms, ``energy'', and so on}

\author{S. Semmes}

\date{}

\maketitle

	John Ball keeps asking questions of the following sort.
Suppose that one has a homeomorphism from a domain in a Euclidean
space onto its image in the same Euclidean space.  Assume also that
the homeomorphism has ``finite energy'' with respect to some
reasonable functional, which would normally entail something like
first distributional derivatives in $L^p$ for some $p$, $1 \le p <
\infty$, and the inverses of the associated differentials lying in some
$L^q$, $1 \le q < \infty$.  The latter might be controlled in terms
of integrability properties of the reciprocal of the Jacobian of
the mapping (the determinant of the differential).  Under conditions
such as these, what kinds of approximations of the homeomorphism
can one make by more regular homeomorphisms, approximations which
respect similar integrability conditions for the differentials and
their inverses?

	Let us restrict ourselves to dimensions $3$ and lower, since
all sorts of strange things happen in dimensions $4$ and larger, and
since dimensions less than or equal to $3$ are physically relevant
(elasticity theory, etc., as Dr.\ Ball well knows).  This type of
issue in dimension $1$ can be treated in a direct manner, by writing
the mapping as the integral of its derivative, and so we focus on
dimensions $2$ and $3$.

	There are very famous results about approximating
homeomorphisms by piecewise-linear homeomorphisms in dimensions $2$
and $3$.  See \cite{Bing, Moise}.  More precisely, these are
approximations in $C^0$ senses, which are already quite nontrivial and
useful in the study of topology.  The results include \emph{relative}
versions, in which a homeomorphism is regularized in some parts while
not changing it on other parts where it is already regular.

	What about approximations which also respect the differential
in some manner?

	A basic strategy in making approximations of a function in a
Sobolev space is to first choose a set on which the function behaves
nicely, and whose complement has small measure.  The nice behavior
might involve boundedness of the first derivatives, continuity of the
first derivatives, and so forth.  Although the complement has small
measure, it typically need not have any simple structure, but could be
quite scatterred and irregular.  The second step would be to modify
the function on the small set.  If there are no additional constraints
on the function, then there are relatively simple techniques of
extension and regularization.  However, if one wants the result to be
a homeomorphism, then this second step becomes much more complicated.

	For more on approximations of homeomorphisms, if not exactly
of this form, see \cite{D-S, Luu3, Sullivan, TV2, V1}.

	This type of conundrum seems a bit odd to me, in that for a
number of basic situations that would arise in a simple way, there
should not be as much trouble.  One way to look at this is that often
there is something like a one-parameter family of mappings.  A related
issue is that in topology a basic point is often to construct
\emph{isotopies} between homeomorphisms, i.e., a continuous family of
homeomorphisms.  This is quite different from \emph{homotopies}, which
are continuous families of mappings which are not required to be
injective, even if the mappings that the homotopy goes between are
injective.

	In particular, there are a number of results to the effect
that homeomorphisms which are close in a $C^0$ sense can be connected
by an isotopy, as in \cite{Cer, E-K, Fisher, Hamstrom, Kister1,
Kister2, Luu1, Sullivan}, and where the isotopy stays close to the
original homeomorphisms.  This is part of the reason that a $C^0$
approximation can be useful, since otherwise it seems rather weak.

	In another direction, let us recall a famous result of Hatcher
\cite{Hatcher1, Hatcher2, Lau} concerning embedded two-dimensional
spheres in ${\bf R}^3$.  More precisely, one considers smoothly embedded
two-dimensional spheres in ${\bf R}^3$, and the space of these can be
locally identified with the space of smooth real-valued functions on the
$2$-sphere, because of the tubular neighborhood theorem.  Hatcher's
result says that this space is \emph{contractable}.

	For the analogous question in the plane, the Riemann mapping
theorem can be used.  Paul Schweitzer keeps asking about possible 
analytic proofs for two-dimensional spheres in ${\bf R}^3$.

	Note that embedded $2$-spheres in ${\bf R}^3$ can be viewed as
the boundaries of solid $3$-dimensional balls, just as Jordan curves
in the plane can be viewed as boundaries of $2$-dimensional disks.

	For another very interesting direction along the lines of
analysis and complexity of shapes, see \cite{C-K-S, Ku-S1, Ku-S2}.

\end{document}